### v1

@@ 4.1.2 Implementation Details
Our DiffMark is implemented by PyTorch [32] and executed on NVIDIA RTX 3090Ti. The pre-trained frozen autoencoder in the training phase is VQGAN [21]. Our DiffMark is trained on CelebA-HQ dataset for 151.2k steps with a batch size of 16, which is equivalent to 100 epochs. We use the AdamW optimizer [33] with a learning rate of 1e-4. In Equation 12, the weight $\alpha$ is set to 0.1. To balance the visual quality and watermark robustness, we initialize the weight $\beta$ to 10 in Equation 15, then reduce it to 1 after 5k and 10k optimization steps for the 128×128 and 256×256 resolutions respectively. For the embedding dimensionality of the embedding table, we set it to 256 for 128×128 images and 1024 for 256×256 images respectively. To conserve memory under limited computation, we precompute the noised-watermarked image gap with detached computation graphs, then adding it back to preserve gradient flow.
@@ 4.1.3 Comparison
The contrastive methods encompass several deep watermarking methods, including MBRS [13], CIN [23], ARWGAN [14], SepMark [10], EditGuard [25] and LampMark [12]. For SepMark [10], we directly adopted the official pre-trained weights, as the dataset and experimental settings are largely consistent. For the other methods, we used their officially released codes and trained them on the CelebA-HQ dataset for 100 epochs. Since our objective was to compare the robustness of binary watermark, we only trained the binary watermark network in EditGuard [25]. Furthermore, we standardized the length of watermark to 30 and 128 for facial images with resolutions of 128×128 and 256×256, respectively. We re-construct landmark perceptual watermarks for LampMark [12], following its original method, to achieve consistent watermark length with other comparative methods. To compare the robustness of watermarks against Deepfake manipulations, we selected SimSwap [1], UniFace [34], CSCS [35], StarGAN [2], and FSRT [3] as typical Deepfake methods, covering three major Deepfake categories: face swapping, face attribute editing and face reenactment.
@@ 4.2.1 Visual Quality
In the watermark embedding invisibility experiment, we evaluated the average PSNR, SSIM and LPIPS between the original and watermarked image. These three metrics are used to assess the differences in image quality, structural similarity, and perceptual similarity between the watermarked image and the original image. As shown in Table 1, our DiffMark maintains the best visual quality at both 128×128 and 256×256 image resolutions, outperforming existing watermarking methods. This indicates that the watermarked images generated by our DiffMark are very similar to the original images, with almost no perceptible visual differences. The performance of ARWGAN [14] substantially deteriorates at 256×256 resolution, primarily due to its difficulties in network convergence at slightly higher resolutions.
@@
MBRS [13] 35.1897 0.9021 0.0744 36.3383 0.8857 0.1188
CIN [23] 39.7044 0.9308 0.0248 37.1549 0.8483 0.0705
ARWGAN [14] 38.5746 0.9733 0.0183 29.6473 0.8271 0.2792
SepMark [10] 38.3129 0.9599 0.0196 38.4669 0.9339 0.0407
EditGuard [25] 37.1664 0.9516 0.0746 36.6557 0.8910 0.1496
LampMark [12] 40.2231 0.9666 0.0293 39.5722 0.9515 0.0715
Ours 41.3005 0.9776 0.0090 41.9620 0.9769 0.0116
@@ 4.2.2 Robustness
MBRS [13] CIN [23] ARWGAN [14] SepMark [10] EditGuard [25] LampMark [12] Ours
@@
Resize(p=0.8) 10.72% 13.58% 24.97% 36.21% 0.02% 10.10% 23.81% 3.40% 0.60% 0.28% 14.00% 13.82% 0.01% 0.01%
Dropout(p=0.6) 0.71% 8.64% 0.00% 0.00% 0.69% 12.44% 0.35% 0.28% 1.02% 0.97% 1.91% 2.43% 0.32% 0.60%
GaussianNoise(s=0.1) 0.07% 0.08% 0.00% 0.00% 20.31% 11.65% 0.76% 0.06% 1.67% 1.02% 9.57% 7.89% 1.30% 2.35%
SaltPepper(p=0.1) 12.49% 12.40% 0.00% 0.00% 0.00% 9.84% 0.02% 0.00% 0.09% 0.12% 24.09% 23.82% 0.09% 0.50%
@@
Hue(f=0.1) 0.00% 0.00% 0.00% 0.00% 2.56% 15.68% 0.54% 0.00% 0.34% 0.17% 0.00% 0.00% 0.25% 0.45%
JpegTest(Q=50) 0.00% 0.01% 5.25% 8.90% 16.47% 15.34% 1.22% 0.10% 1.42% 3.50% 0.69% 1.83% 0.85% 1.86%
Average 2.48% 4.22% 3.16% 3.85% 4.19% 13.43% 2.28% 0.33% 1.13% 1.37% 4.53% 4.31% 0.30% 0.60%
In Table 2, we evaluated the method using various distortions such as {Identity, Resize, Dropout, GaussianNoise, SaltPepper, GaussianBlur, MedianBlur, Brightness, Contrast, Saturation, Hue, JpegTest}. It can be seen that our DiffMark, along with the SepMark, achieves the lowest average watermark BER at 128×128 and 256×256 resolutions respectively. However, since our DiffMark generates watermarked images by denoising based on the standard Gaussian distribution, the BER slightly increases when facing Gaussian noise, indicating some sensitivity to this type of distortion. It is observed that the compared watermarking methods consistently exhibit vulnerability to specific distortions like resize and salt-and-pepper noise, likely due to insufficient consideration of these distortions in their network architecture or noise layer design. Notably, during training, we only incorporated a frozen-parameter autoencoder and did not include these common distortions in our end-to-end training framework. Despite this, our DiffMark still maintains strong robustness against these distortions, which exceeds our expectations.
@@
MBRS [13] CIN [23] ARWGAN [14] SepMark [10] EditGuard [25] LampMark [12] Ours
@@
SimSwap [1] 24.60% 27.09% 40.08% 31.07% 46.60% 41.57% 20.02% 11.95% 45.32% 47.99% 16.53% 15.11% 5.58% 5.98%
UniFace [34] 0.48% 26.57% 11.80% 48.27% 26.28% 42.69% 0.34% 31.93% 9.17% 49.41% 6.28% 29.85% 0.01% 2.17%
CSCS [35] 10.10% 10.33% 0.29% 0.79% 6.29% 14.82% 0.68% 2.35% 0.99% 1.61% 2.30% 0.63% 0.13% 0.54%
StarGAN [2] 5.49% 17.26% 56.93% 38.56% 36.78% 32.35% 0.11% 0.01% 7.62% 2.12% 7.30% 4.05% 4.63% 3.83%
FSRT [3] 2.40% 20.92% 3.20% 35.22% 4.34% 35.32% 0.78% 9.21% 5.77% 31.17% 2.93% 18.08% 0.12% 4.09%
VQGAN [21] 0.05% 0.73% 39.60% 24.90% 35.06% 34.86% 1.28% 0.10% 7.49% 6.38% 10.22% 10.75% 0.02% 0.01%
@@
In Table 3, we conducted the evaluation of BER under representative Deepfake manipulations such as {SimSwap, UniFace, CSCS, StarGAN, FSRT}. In this study, the VQGAN serves as an autoencoder to simulate Deepfake manipulations in our training framework and we include it in Table 3 and Table 4 for comparative reference. The experimental results demonstrate that our method achieves lower BER against most Deepfake manipulations, outperforming many of comparison methods. Notably, SepMark exhibits the lowest BER on StarGAN, which may be due to the targeted optimization of StarGAN within its training framework. However, when considering the average BER across both 128×128 and 256×256 resolutions, our DiffMark achieves the best performance, confirming its generalization capability in handling various Deepfake manipulations. Furthermore, we observe that even the same Deepfake model can exhibit varying impacts on watermark robustness across different image resolutions, as seen with UniFace [34] and FSRT [3]. It may be because the Deepfake model tends to bring more distortions to the watermarked image as the image resolution increases.
@@ 4.3 Cross-Dataset Evaluation
To further evaluate the generalization of DiffMark in cross-dataset settings, we conducted experiments on the LFW dataset at both 128×128 and 256×256 resolutions.
@@
MBRS [13] CIN [23] ARWGAN [14] SepMark [10] EditGuard [25] LampMark [12] Ours
@@
SimSwap [1] 26.38% 24.23% 43.02% 27.47% 47.74% 42.56% 25.91% 18.07% 45.53% 47.40% 24.95% 15.61% 8.81% 8.37%
UniFace [34] 0.14% 21.16% 12.26% 47.99% 26.90% 42.34% 0.41% 25.52% 8.98% 48.80% 10.17% 31.00% 0.01% 2.10%
CSCS [35] 5.27% 6.53% 0.31% 0.70% 14.62% 15.12% 1.73% 1.06% 4.29% 2.33% 10.32% 1.38% 0.65% 3.85%
StarGAN [2] 6.19% 17.29% 58.15% 43.18% 40.66% 33.29% 0.51% 0.04% 12.94% 3.01% 17.00% 5.49% 5.54% 5.43%
FSRT [3] 4.01% 19.69% 7.84% 38.36% 11.64% 36.83% 1.54% 14.26% 12.90% 35.65% 15.21% 25.46% 0.49% 10.07%
VQGAN [21] 0.17% 0.13% 40.13% 18.10% 39.82% 33.28% 1.17% 0.16% 14.38% 5.87% 21.53% 6.49% 0.03% 0.02%
Average 7.03% 14.84% 26.95% 29.30% 30.23% 33.90% 5.21% 9.85% 16.50% 23.84% 16.53% 14.24% 2.59% 4.97%
PSNR ↑ 34.99 36.73 39.64 37.21 38.60 29.69 37.30 38.28 34.98 34.23 37.18 39.56 39.08 41.33
SSIM ↑ 0.900 0.880 0.930 0.824 0.973 0.851 0.951 0.930 0.938 0.840 0.937 0.947 0.969 0.973
LPIPS ↓ 0.080 0.144 0.025 0.094 0.019 0.178 0.027 0.055 0.093 0.217 0.056 0.085 0.014 0.011

### v2

@@ 4.1.2 Implementation Details
Our DiffMark is implemented by PyTorch Paszke et al. [2019] and executed on NVIDIA RTX 3090Ti. The pre-trained frozen autoencoder in the training phase is VQGAN Esser et al. [2021]. Our DiffMark is trained on CelebA-HQ dataset for 151.2k steps with a batch size of 16, which is equivalent to 100 epochs. We use the AdamW optimizer Loshchilov and Hutter [2017] with a learning rate of 1e-4. In Equation 12, the weight $\alpha$ is set to 0.1. To balance the visual quality and watermark robustness, we initialize the weight $\beta$ to 10 in Equation 15, then reduce it to 1 after 5k and 10k optimization steps for the $128\times 128$ and $256\times 256$ resolutions respectively. For the embedding dimensionality of the embedding table, we set it to 256 for $128\times 128$ images and 1024 for $256\times 256$ images respectively. To conserve memory under limited computation, we precompute the noised-watermarked image gap with detached computation graphs, and then add it back to preserve gradient flow. Considering the long sampling time with 1000 steps in the original diffusion model, we reduced the training diffusion steps to 100 and used the DDIM sampler Song et al. [2020] with 10 steps during sampling.
@@ 4.1.3 Comparison
The contrastive methods encompass several deep watermarking methods, including MBRS Jia et al. [2021], CIN Ma et al. [2022], ARWGAN Huang et al. [2023], SepMark Wu et al. [2023], EditGuard Zhang et al. [2024] and LampMark Wang et al. [2024]. For SepMark Wu et al. [2023], we directly adopted the official pre-trained weights, as the dataset and experimental settings are largely consistent. For the other methods, we used their officially released codes and trained them on the CelebA-HQ dataset for 100 epochs. Since our objective was to compare the robustness of binary watermark, we only trained the binary watermark network in EditGuard Zhang et al. [2024]. Furthermore, we standardized the length of watermark to 30 and 128 for facial images with resolutions of $128\times 128$ and $256\times 256$, respectively. We re-construct landmark perceptual watermarks for LampMark Wang et al. [2024], following its original method, to achieve consistent watermark length with other comparative methods. To compare the robustness of watermarks against Deepfake manipulations, we selected SimSwap Chen et al. [2020], UniFace Xu et al. [2022], CSCS Huang et al. [2024], StarGAN Choi et al. [2018], and FSRT Rochow et al. [2024] as typical Deepfake methods, covering three major Deepfake categories: face swapping, face attribute editing and face reenactment.
@@ 4.2.1 Visual Quality
In the watermark embedding invisibility experiment, we evaluated the average PSNR, SSIM and LPIPS between the original and watermarked image. These three metrics are used to assess the differences in image quality, structural similarity, and perceptual similarity between the watermarked image and the original image. As shown in Table 1, our DiffMark maintains the best visual quality at both $128\times 128$ and $256\times 256$ image resolutions, outperforming existing watermarking methods. This indicates that the watermarked images generated by our DiffMark are very similar to the original images, with almost no perceptible visual differences. The performance of ARWGAN Huang et al. [2023] substantially deteriorates at $256\times 256$ resolution, primarily due to its difficulties in network convergence at slightly higher resolutions.
@@
MBRS Jia et al. [2021] 35.1897 0.9021 0.0744 36.3383 0.8857 0.1188
CIN Ma et al. [2022] 39.7044 0.9308 0.0248 37.1549 0.8483 0.0705
ARWGAN Huang et al. [2023] 38.5746 0.9733 0.0183 29.6473 0.8271 0.2792
SepMark Wu et al. [2023] 38.3129 0.9599 0.0196 38.4669 0.9339 0.0407
EditGuard Zhang et al. [2024] 37.1664 0.9516 0.0746 36.6557 0.8910 0.1496
LampMark Wang et al. [2024] 40.2231 0.9666 0.0293 39.5722 0.9515 0.0715
Ours 41.2869 0.9776 0.0090 41.9572 0.9769 0.0116
@@ 4.2.2 Robustness
MBRS Jia et al. [2021] CIN Ma et al. [2022] ARWGAN Huang et al. [2023] SepMark Wu et al. [2023] EditGuard Zhang et al. [2024] LampMark Wang et al. [2024] Ours
@@
Resize(p=0.8) 10.72% 13.58% 24.97% 36.21% 0.02% 10.10% 23.81% 3.40% 0.60% 0.28% 14.00% 13.82% 0.02% 0.01%
Dropout(p=0.6) 0.71% 8.64% 0.00% 0.00% 0.69% 12.44% 0.35% 0.28% 1.02% 0.97% 1.91% 2.43% 0.31% 0.60%
GaussianNoise(s=0.1) 0.07% 0.08% 0.00% 0.00% 20.31% 11.65% 0.76% 0.06% 1.67% 1.02% 9.57% 7.89% 1.35% 2.37%
SaltPepper(p=0.1) 12.49% 12.40% 0.00% 0.00% 0.00% 9.84% 0.02% 0.00% 0.09% 0.12% 24.09% 23.82% 0.09% 0.48%
@@
Hue(f=0.1) 0.00% 0.00% 0.00% 0.00% 2.56% 15.68% 0.54% 0.00% 0.34% 0.17% 0.00% 0.00% 0.27% 0.42%
JpegTest(Q=50) 0.00% 0.01% 5.25% 8.90% 16.47% 15.34% 1.22% 0.10% 1.42% 3.50% 0.69% 1.83% 0.85% 1.81%
Average 2.48% 4.22% 3.16% 3.85% 4.19% 13.43% 2.28% 0.33% 1.13% 1.37% 4.53% 4.31% 0.30% 0.59%
In Table 2, we evaluated the method using various distortions such as {Identity, Resize, Dropout, GaussianNoise, SaltPepper, GaussianBlur, MedianBlur, Brightness, Contrast, Saturation, Hue, JpegTest}. It can be seen that our DiffMark, along with the SepMark, achieves the lowest average watermark BER at $128\times 128$ and $256\times 256$ resolutions respectively. However, since our DiffMark generates watermarked images by denoising based on the standard Gaussian distribution, the BER slightly increases when facing Gaussian noise, indicating some sensitivity to this type of distortion. It is observed that the compared watermarking methods consistently exhibit vulnerability to specific distortions like resize and salt-and-pepper noise, likely due to insufficient consideration of these distortions in their network architecture or noise layer design. Notably, during training, we only incorporated a frozen-parameter autoencoder and did not include these common distortions in our end-to-end training framework. Despite this, our DiffMark still maintains strong robustness against these distortions, which exceeds our expectations.
@@
MBRS Jia et al. [2021] CIN Ma et al. [2022] ARWGAN Huang et al. [2023] SepMark Wu et al. [2023] EditGuard Zhang et al. [2024] LampMark Wang et al. [2024] Ours
@@
SimSwap Chen et al. [2020] 24.60% 27.09% 40.08% 31.07% 46.60% 41.57% 20.02% 11.95% 45.32% 47.99% 16.53% 15.11% 5.58% 5.96%
UniFace Xu et al. [2022] 0.48% 26.57% 11.80% 48.27% 26.28% 42.69% 0.34% 31.93% 9.17% 49.41% 6.28% 29.85% 0.01% 2.20%
CSCS Huang et al. [2024] 10.10% 10.33% 0.29% 0.79% 6.29% 14.82% 0.68% 2.35% 0.99% 1.61% 2.30% 0.63% 0.13% 0.56%
StarGAN Choi et al. [2018] 5.49% 17.26% 56.93% 38.56% 36.78% 32.35% 0.11% 0.01% 7.62% 2.12% 7.30% 4.05% 4.66% 3.82%
FSRT Rochow et al. [2024] 2.40% 20.92% 3.20% 35.22% 4.34% 35.32% 0.78% 9.21% 5.77% 31.17% 2.93% 18.08% 0.12% 4.05%
VQGAN Esser et al. [2021] 0.05% 0.73% 39.60% 24.90% 35.06% 34.86% 1.28% 0.10% 7.49% 6.38% 10.22% 10.75% 0.02% 0.02%
@@
In Table 3, we conducted the evaluation of BER under representative Deepfake manipulations such as {SimSwap, UniFace, CSCS, StarGAN, FSRT}. In this study, the VQGAN serves as an autoencoder to simulate Deepfake manipulations in our training framework and we include it in Table 3 and Table 4 for comparative reference. The experimental results demonstrate that our method achieves lower BER against most Deepfake manipulations, outperforming many of comparison methods. Notably, SepMark exhibits the lowest BER on StarGAN, which may be due to the targeted optimization of StarGAN within its training framework. However, when considering the average BER across both $128\times 128$ and $256\times 256$ resolutions, our DiffMark achieves the best performance, confirming its generalization capability in handling various Deepfake manipulations. Furthermore, we observe that even the same Deepfake model can exhibit varying impacts on watermark robustness across different image resolutions, as seen with UniFace Xu et al. [2022] and FSRT Rochow et al. [2024]. It may be because the Deepfake model tends to bring more distortions to the watermarked image as the image resolution increases.
@@ 4.3 Cross-Dataset Evaluation
To further evaluate the generalization of DiffMark in cross-dataset settings, we conducted experiments on the LFW dataset at both $128\times 128$ and $256\times 256$ resolutions.
@@
MBRS Jia et al. [2021] CIN Ma et al. [2022] ARWGAN Huang et al. [2023] SepMark Wu et al. [2023] EditGuard Zhang et al. [2024] LampMark Wang et al. [2024] Ours
@@
SimSwap Chen et al. [2020] 26.38% 24.23% 43.02% 27.47% 47.74% 42.56% 25.91% 18.07% 45.53% 47.40% 24.95% 15.61% 9.03% 8.35%
UniFace Xu et al. [2022] 0.14% 21.16% 12.26% 47.99% 26.90% 42.34% 0.41% 25.52% 8.98% 48.80% 10.17% 31.00% 0.01% 2.07%
CSCS Huang et al. [2024] 5.27% 6.53% 0.31% 0.70% 14.62% 15.12% 1.73% 1.06% 4.29% 2.33% 10.32% 1.38% 0.64% 3.82%
StarGAN Choi et al. [2018] 6.19% 17.29% 58.15% 43.18% 40.66% 33.29% 0.51% 0.04% 12.94% 3.01% 17.00% 5.49% 6.03% 5.45%
FSRT Rochow et al. [2024] 4.01% 19.69% 7.84% 38.36% 11.64% 36.83% 1.54% 14.26% 12.90% 35.65% 15.21% 25.46% 0.43% 9.93%
VQGAN Esser et al. [2021] 0.17% 0.13% 40.13% 18.10% 39.82% 33.28% 1.17% 0.16% 14.38% 5.87% 21.53% 6.49% 0.04% 0.01%
Average 7.03% 14.84% 26.95% 29.30% 30.23% 33.90% 5.21% 9.85% 16.50% 23.84% 16.53% 14.24% 2.70% 4.94%
PSNR ↑ 34.99 36.73 39.64 37.21 38.60 29.69 37.30 38.28 34.98 34.23 37.18 39.56 39.10 41.33
SSIM ↑ 0.900 0.880 0.930 0.824 0.973 0.851 0.951 0.930 0.938 0.840 0.937 0.947 0.970 0.973
LPIPS ↓ 0.080 0.144 0.025 0.094 0.019 0.178 0.027 0.055 0.093 0.217 0.056 0.085 0.013 0.011
@@
4.4 Ablation Studies
4.4.1 Hyperparameters $\alpha$ and $\beta$
Table 5: Ablation Study on the Hyperparameters $\alpha$ and $\beta$ (Image Size $128\times 128$). The Table Shows Image Quality Metrics (PSNR, SSIM, LPIPS) and Watermark Bit Error Rates (BER) under Various Deepfake Manipulations.
$\alpha$ $\beta$ Image Quality Metrics Watermark Bit Error Rates (BER) $\downarrow$
[…]
0.1 1.0 37.9254 0.9579 0.0202 2.86% 0.00% 0.06% 2.60% 0.01% 0.00% 0.92%
0.1 0.1 48.1395 0.9946 0.0007 50.14% 49.93% 49.80% 50.43% 50.35% 50.09% 50.12%
0.1 $1.0\to 0.1$ 41.2869 0.9776 0.0090 5.58% 0.01% 0.13% 4.66% 0.12% 0.02% 1.75%
1.0 $1.0\to 0.1$ 40.1124 0.9657 0.0047 9.50% 0.01% 0.17% 10.39% 0.66% 0.16% 3.48%
0.01 $1.0\to 0.1$ 41.2220 0.9743 0.0165 2.69% 0.00% 0.01% 2.03% 0.07% 0.00% 0.80%
In this section, we perform an ablation study on two hyperparameters, $\alpha$ and $\beta$, to evaluate their effects on the watermarking performance. The results in Table 5 highlight the trade-off between watermark robustness and image quality. When $\alpha$ is fixed at 0.1, the model fails to converge with $\beta=0.1$. Increasing $\beta$ to 1.0 reduces the average bit error rate (BER) but at the cost of lower image quality. Annealing $\beta$ from 1.0 to 0.1 strikes a better balance between robustness and invisibility. Similarly, with $\beta$ annealed from 1.0 to 0.1, a smaller $\alpha$ value further decreases BER, yet also degrades image quality. However, increasing $\alpha$ to 1.0 significantly diminishes watermark robustness without improving image quality. In comparison, setting $\alpha=0.1$ yields a more desirable outcome in terms of both metrics. These results indicate that improved robustness often comes at the expense of visual fidelity—smaller $\alpha$ and larger $\beta$ could enhance robustness but reduce image quality. The optimal balance is dictated by application priorities.
4.4.2 Cross Information Fusion Module
To analyze the impact of the embedding dimension in the Cross Information Fusion (CIF) module, we conduct an ablation study on the dimension of the embedding table.
Table 6: Ablation Study on the Dimension of Embedding Table in the CIF Module (Image Size $128\times 128$). We Report Image Quality (PSNR, SSIM, LPIPS) and Watermark Bit Error Rates (BER) under Various Deepfake Manipulations.
Embedding Dim Image Quality Metrics Watermark Bit Error Rates (BER) $\downarrow$
[…]
128 41.1802 0.9754 0.0104 6.71% 0.02% 0.31% 5.89% 0.12% 0.03% 2.18%
256 41.2869 0.9776 0.0090 5.58% 0.01% 0.13% 4.66% 0.12% 0.02% 1.75%
512 40.8905 0.9744 0.0104 5.09% 0.00% 0.07% 3.70% 0.18% 0.02% 1.51%
1024 40.7391 0.9749 0.0107 4.39% 0.00% 0.04% 3.18% 0.16% 0.02% 1.30%
1536 40.3886 0.9716 0.0118 5.16% 0.00% 0.03% 3.53% 0.20% 0.02% 1.49%
Table 6 shows that as the embedding dimension increases, the average bit error rate (BER) generally decreases, indicating improved robustness against Deepfake manipulations. However, this improvement plateaus beyond 1024 dimensions, as seen by the rise in BER at 1536 dimensions. This suggests that while increasing the embedding dimension initially enhances robustness, the gains diminish beyond a certain point, and further increases contribute little to additional improvement in watermark robustness. Moreover, the enhanced robustness is often associated with a decrease in image quality, as reflected by lower PSNR and SSIM values and higher LPIPS values.